40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 45 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  14853   Thu Aug 22 20:56:51 2019 KojiUpdateCDSMC1 glitch removed (for now) and IMC locking recovered

The internal ribbon cable for the MC1 satellite box was replaced with the one in the spare box. The MC1 box was closed and reinstalled as before. The IMC is locking well.

Now the burnt cable was disassembled and reassembles with a new cable. It is now in the spare box.

The case closed (literally)

  14852   Thu Aug 22 12:54:06 2019 KojiUpdateCDSMC1 glitch removed (for now) and IMC locking recovered

I have checked the MC1 satellite box and made a bunch of changes. For now, the glitches coming from the satellite box is gone. I quickly tested the MC1 damping and the IMC locking. The IMC was locked as usual. I still have some cleaning up but will work on them today and tomorrow.


Attachment 1: Result

The noise level of the satellite box was tested with the suspension simulator (i.e., five pair of the LED and PD in a plastic box).

Each plot shows the ASD of the sensor outputs 1) before the modification, 2) after the change, and 3) with the satellite box disconnected (i.e., the noise from the PD whitening filter in the SUS rack).

Before the modification, these five signals showed significant (~0.9) correlation each other, indicating that the noise source is common. After the modification, the spectra are lowered down to the noise level of the whitening filters, and there is no correlation observed anymore. EXCEPT FOR the LR sensor: It seems that the LR has additional noise issue somewhere in the downstream. This is a separate issue.


Attachment 2: Photo of the satellite box before the modification

The thermal environment in the box is terrible. They are too hot to touch. You can see that the flat ribbon cable was burned. The amps, buffers, and regulators generate much heat.

Attachment 3: Where the board was modified

- (upper left corner) Every time I touched C51, the diode output went to zero. So C51 was replaced with WIMA 10uF (50V) cap.

- (lower left area) I found a clear indication of the glitch coming from the PD bias path (U3C). So I first replaced another 10uF (C50) with WIMA 10uF (50V). This did not change the glitch. So I replaced U3 (LT1125). This U3 had unused opamp which had railed to the supply voltage. Pins 14 and 15 of U3 were shorted to ground.

- (lower right corner) Similarly to U3, U6 also had two opamps which are railed due to no termination. U6 was replaced, and Pins 11, 12, 14, and 15 were shorted to ground.

- (middle right) During the course of the search, I suspected that the LR glitch comes from U5. So U5 was replaced to the new chip, but this had no effect.


Attachment 4: Thermal degradation of the internal ribbon cable

Because of the heat, the internal ribbon cable lost the flexibility. The cable is cracked and brittle. It now exposes  some wires. This needs to be replaced. I'll work on this later this week.

Attachment 5: Thermal degradation of the board

Because of the excessive heat for those 20years, the bond between the board and the patten were degraded. In conjunction with extremely thin wire pattern, desoldering of the components (particularly LT1125s) was very difficult. I'd want to throw away this board right now if it were possible...

Attachment 6: Shorting the unused opamps

This shows how the pieces of wires were soldered to ground vias to short the unused opamps.


Attachment 7: Comparison of the noise level with the sus simulator and the actual MC1 motion

After the satellite box fix, the sensor outputs were measured with the suspension connected. This shows that the suspension is moving much more than the noise level around 1Hz. However, at the microseismic frequency there is also most no mergin. Considering the use of the adaptive feedforward, we need to lower the noise of the satellite box as well as the noise of the whitening filters.

=> Use better chips (no LT1125, no current buffers), use low noise resistors, better thermal environment.


Attachment 1: satellite_box.pdf
satellite_box.pdf
Attachment 2: before.jpg
before.jpg
Attachment 3: after.jpg
after.jpg
Attachment 4: P_20190821_194035.jpg
P_20190821_194035.jpg
Attachment 5: P_20190821_174240.jpg
P_20190821_174240.jpg
Attachment 6: P_20190821_194013.jpg
P_20190821_194013.jpg
Attachment 7: comparison_satellite_box.pdf
comparison_satellite_box.pdf
  14851   Tue Aug 20 19:05:24 2019 KojiUpdateCDSMC1 (and MC3) troubleshoot

Started the troubleshoot from the MC1 issue. Gautam showed me how to use the fake PD/LED pair to diagnose the satellite box without involving the suspension mechanics.

This revealed that the MC1 has frequent light level glitches which are common for five sensors. This feature does not exist in the test with the MC3 satellite box. I will open and check the MC1 satellite box to find the cause of this common glitches tomorrow. MC1 is currently shutdown and undamped.

BTW, at the MC3 test, i found that J2 of the satellite box (male Dsub) has all the pins too low (or too short?). I brought the box outside and found that the housing of this connector was half broken down. The connector was reassembled and the metal parts of the housing was bent again so that the housing can hold the connector body tightly.

The MC3 satellite box was restored and connected to the cables. As I touched this box, it is still under probation.

Attachment 1: Screenshot_from_2019-08-20_17-26-01.png
Screenshot_from_2019-08-20_17-26-01.png
Attachment 2: Screenshot_from_2019-08-20_17-43-03.png
Screenshot_from_2019-08-20_17-43-03.png
  14850   Mon Aug 19 14:36:21 2019 gautamUpdateCDSc1iscaux remaining work

Here is what is left to do:

  1. Strain relief of all cabling. Chub will take care of this in the coming days. I have said he can connect and disconnect cables as he pleases, but after this work, we may require a hard reboot of the Acromag chassis before restoring functionality to the channels, as it is known that the Acromags can sometimes get "stuck" by a sudden connection of voltage.
  2. Installation of DB15 cable to the P2 connector of the CM board and a DB9 cable to the ALS demod unit (LO and RF power monitors). These will arrive in the next couple of days and Chub will take care of the install.
  3. Design, manufacture and install of a custom version of the backplane P1 adaptor board with only 1 D37 connector - for some of the PD DC signals, a custom adaptor board, part number D010005 for which I can't find any schematics is already installed on the P2 connector, and makes the DC monitor signals available to 4 LEMO connectors. These signals are then digitized by the fast CDS system, presumably for PDH signal normalization. The footprint of this P2--->LEMO adaptor is such that we cannot simply install our P1---> 2xDB37 adaptor boards, because of space constraints. Fortunately, there is a simple fix to reduce the footprint of the board: remove the bottom DB37 connector, which is unused in the c1iscaux system except for the CM board. I recommend getting ~10 pcs of such boards, as it is also useful in a few other places, where the power cabling to the eurocrates are a space constraint. See Attachment #1 for a picture explaining this situation. Anyone want to volunteer to take care of this?
  4. In-situ testing. This is easiest done with some light available in the interferometer. Which in turn requires IMC to be locked. Which in turn requires satellite box fixing. Anyone want to volunteer to take care of this?
  5. Modify C0EDCU.ini to trend the new slow channels we may want long-term monitoring of (e.g. LO power levels to the Demod boards). Anyone want to volunteer to take care of this?
  6. Decide what to do about the CM latch logic. There are some contraints with the way the acromag register addressing works, that I've had to change the way the mbboDirect bits are controlled. Unfortunately, this seems to sometimes and unpredictably cause the bits to flip in a non-robust way, which is the whole point of having the latch in the first place. Either the latch logic needs to be improved, or we need to implement the latch logic in the fast CDS system, not the slow.

Today I set up the autoburt.req file for the c1iscaux channels, and confirmed that the snapshots are getting recorded. There were a lot of channels in the old autoburt.req file which I thought were un-necessary (and several which no longer exist), so now the only channels that are burt-ed are the whitening gains and states of the AA filters. If someone feels we need more channels to be snapshot recorded, you can add them to the file.

In the old target directory, there were also various versions of a "saverestore.req" file - why do we need this in addition to an autoburt? I guess it is possible they are used by the IFOconfigure scripts to setup some whitening gains etc...

Attachment 1: caseForSmallerFootprint.pdf
caseForSmallerFootprint.pdf
  14849   Sat Aug 17 16:49:23 2019 gautamUpdateCDSMore 1Y3 work

Work done today:

  1. All ribbon cable connections to the backplane of the 1Y2 Eurocrates were removed. The cables themselves were cleared for more space to work with.
  2. 20x 15ft DB37 Cables were run between 1Y2 and 1Y3 via overhead cable tray.
  3. Backplane interface boards were installed for 1Y2 Eurocrate boards.
  4. Connections were made between the Acromag chassis and the eurocrate electronics modules.

Testing of functionality:

  1. Fast BIO switching was verified to work for the following photodiodes:
    • AS55, AS110, REFL11, REFL33, REFL55, REFL165, POX11, POY11, POP22, POP110.
    • No light was incident on the PDs.
    • Test was done by increasing the whitening gain to +45 dB, and then looking at the ASD of the electronics noise between 50 Hz and 500 Hz with the whitening enabled/disabled. We expect x10 difference between the two states. This was seen.
  2. "DetMon" channels were verified to work - see Attachment #1
    • Y-axis units is volts
    • Test was done by toggling the output of the 11 MHz Marconi, and looking for a change.
    • As seen in the attachment, all 5 monitor channels show a change.
    • This needs to be calibrated into some sensible units - I don't know why the different modulation frequencies have such different readbacks from supposedly identical Demod Board monitor points.
    • Not sure if the ~10 V reported by the REFL165 monitor point is real or saturated.
    • These channels are installed to signal/help debug the infamous ERA-5 decay problem, but maybe already some are decayed?
  3. QPD interface channels were verified to work - see Attachment #2
    • Test was done by shining a green laser pointer on QPD quadrants.

Much testing remains to be done, but I defer further testing till Monday - the main functionality to be verified in the short run is the whitening gain stepping. The strain-relief of cables and general cleanup will be undertaken by Chub. Current state of affairs is in Attachment #3, leaves much to be desired in terms of cleanliness.

I will also setup the autoburt for the new machine on Monday. We will also need to add some channels to C0EDCU.ini if we want to trend them over some years (e.g. RF signal powers for monitoring ERA-5 health).

* c1lsc FE was rebooted using the usual script, and everything seems to be healthy in CDS-land again, see Attachment #4.

Quote:

Next steps: 

  1. We did not get around to running the DB37 cables between the Acromag chassis and the 1Y2 Eurocrates today - this operation itself took the whole day as we also needed to lay out some support struts etc on the rack to support the Sorensens and the Acromag chassis.
  2. Once the Acromags are connected to the Eurocrates, we have to run in-situ tests to make sure the appropriate functionality has been restored.
  3. We must have bumped something in the c1lsc expansion chassis - the CDS FE overview screen is reporting some errors (see Attachment #3). I will fix this.
  4. General tidiness, strain-relief etc.
Attachment 1: Screen_Shot_2019-08-17_at_3.00.57_PM.png
Screen_Shot_2019-08-17_at_3.00.57_PM.png
Attachment 2: Screen_Shot_2019-08-17_at_3.12.23_PM.png
Screen_Shot_2019-08-17_at_3.12.23_PM.png
Attachment 3: IMG_7804.JPG
IMG_7804.JPG
Attachment 4: Screenshot_from_2019-08-17_17-04-47.png
Screenshot_from_2019-08-17_17-04-47.png
  14848   Fri Aug 16 16:40:04 2019 gautamUpdateCDS1Y3 work

[chub, gautam]

Installation: The following equipment were installed in 1Y3, see Attachment #1:

  1. Supermicro server, which is the new c1iscaux machine, with IP Address 192.168.113.83.
  2. 6U Acromag chassis which contains all the ADCs, DACs and BIO units.
  3. 2 Sorensen DC power supplies to provide +24 V DC and +15 V DC to the Acromags.
  4. Fusable DIN rail power blocks were installed on the North side of the 1Y3 rack - I placed 2 banks of 5 connectors each for +15 V DC and +24 V DC.

Removal: The following equipment was removed from 1Y3:

  1. VME crates that were the old c1iscaux and c1iscaux2 machines.
  2. Spare VME crate that used to be c1susaux, which Chub and I brought over to 1Y3 in an attempt to revive the broken c1iscaux2.
  3. Approximately 30 twisted ribbon cables that were going to the cross connects. For now, we have not done a full cleanup and they are just piled along the east arm (see Attachment #2), beware if you are walking there!

Software: 

  1. I connected the c1iscaux machine to the martian network.
  2. Then I edited the relevant files on chiara to free up the IP addresses previously used by c1iscaux (192.168.113.81) and c1iscaux2 (192.168.113.82), and re-assigned the IP address used for c1iscaux to be 192.168.113.83.
  3. I also changed the hostname of the c1iscaux machine (it was temporarily called c1iscaux3 to allow bench testing).
  4. I moved the old /cvs/cds/caltech/target/c1iscaux and /cvs/cds/caltech/target/c1iscaux2 directories to /cvs/cds/caltech/target/preAcromag_oldVME/c1iscaux and /cvs/cds/caltech/target/preAcromag_oldVME/c1iscaux2 respectively.
  5. I moved the temporarily named /cvs/cds/caltech/target/c1iscaux3 directory, from which I was running all the tests, to /cvs/cds/caltech/target/c1iscaux.
  6. I edited all references to c1iscaux3 in the systemd files so that we can run the approriate systemd services.

Next steps: 

  1. We did not get around to running the DB37 cables between the Acromag chassis and the 1Y2 Eurocrates today - this operation itself took the whole day as we also needed to lay out some support struts etc on the rack to support the Sorensens and the Acromag chassis.
  2. Once the Acromags are connected to the Eurocrates, we have to run in-situ tests to make sure the appropriate functionality has been restored.
  3. We must have bumped something in the c1lsc expansion chassis - the CDS FE overview screen is reporting some errors (see Attachment #3). I will fix this.
  4. General tidiness, strain-relief etc.
Quote:

I judge that we are good to go ahead with an install tomorrow.

Attachment 1: newLook1Y3.JPG
newLook1Y3.JPG
Attachment 2: IMG_7803.JPG
IMG_7803.JPG
Attachment 3: c1lsc_crashed.png
c1lsc_crashed.png
  14847   Fri Aug 16 04:24:03 2019 ranaUpdateALSALS sensing noise due to IMC

What about just use high gain feedback to MC2 below 20 Hz for the IMC lock? That would reduce the excess if this theory is correct.

  14846   Thu Aug 15 18:54:54 2019 gautamUpdateALSALS sensing noise due to IMC

Summary:

I came aross an interesting suggestion by Yutaro that KAGRA's low-frequency ALS noise could be limited by the fact that the IMC comes between the point where the frequencies of the PSL and AUX lasers are sensed (i.e. the ALS beat note), and the point where we want them to be equal (i.e. the input of the arm cavity). I wanted to see if the same effect could be at play in the 40m ALS system. A first estimate suggests to me that the numbers are definitely in the ballpark. If this is true, we may benefit from lower noise ALS by picking off the PSL beam for the ALS beat note after the IMC.

Details:

Even though the KAGRA phase lock scheme is different from the 40m scheme, the algebra holds. I needed an estimate of how much the arm cavity moves, I used data from a POX lock to estimate this. The estimate is probably not very accurate (since the arm cavity length is more stable than the IMC length, and the measured ALS noise, e.g. this elog, is actually better than what this calculation would have me believe), but should be the right order of magnitude. From this crude estimate, it does look like for f<10 Hz, this effect could be significant. I assumed an IMC pole of 3.8 kHz for this calculation.

I've indicated a "target" ALS performance where the ALS noise would be less than the CARM linewidth, which would hopefully make the locking much easier. Seems like realizing this target will be touch-and-go. But if we can implement length feedforward control for the arm cavities using seismometers, the low frequency motion of the optics should go down. It would be interesting to see if the ALS noise gets better at low frequencies with length feedforward engaged.

* Some updates were made to the plot:

  1. Took data from Kiwamu's paper for the seismic noise
  2. Overlaid measured ALS noise
Attachment 1: ALSsensingNoise.pdf
ALSsensingNoise.pdf
  14845   Tue Aug 13 14:36:17 2019 gautamUpdateCDSP1--->P2

As it turns out, only one extra shroud needed to be installed - I did this and migrated the cables for the 4 whitening boards from the P1 to P2 connectors. So until the new Acromag box is installed, we have no control over the whitening gains (slow channels), but do still have control over the whitening filter enable/disable (controlled by fast BIO). I am thinking about the easiest way to test the latter - I think the ambient PD dark noise level is too low to be seen above ADC noise even with the whitening enabled, and setting up drive signals to individual channels is too painful - maybe with +45dB of whitening gain, the (z,p) whitening filter shape can be seen with just PD/demod chain electroncis noise.

Quote:

This morning, I wanted to move the existing cables going to the P1 connectors of the iLIGO whitening boards to the P2 connector, to test the modifications made to allow whitening stage switching. Unfortunately, I found that the shrouds werent installed. Where can I find these?

  14844   Tue Aug 13 08:07:09 2019 gautamUpdateCDSP1--->P2

This morning, I wanted to move the existing cables going to the P1 connectors of the iLIGO whitening boards to the P2 connector, to test the modifications made to allow whitening stage switching. Unfortunately, I found that the shrouds werent installed. Where can I find these?

  14843   Mon Aug 12 21:25:19 2019 KojiUpdateCDSMore bench test of c1iscaux

1.

> Looking through the manual, I found a recommendation (pg10) that the "IN-" terminal of the Acromag ADC units be tied to the "RTN" pins on the same units. I don't know if this preserves the differential receiving capability of the Acromag ADCs

I suppose, we loose the differential capability of an input if the -IN is connected to whatever defined potential. We should check if the channels are still working as a true differential or not.

 

2. If the multi bit operation is too complicated to solve, we can use EPICS Calc channels to breakout a value to bits and send the individual bits as same as the other individual binary channels.

 

  14842   Mon Aug 12 19:58:23 2019 gautamUpdateIOOMC1 suspension oddness

Repair plan:

  1. Get "spare" satellite box working --- Chub
    • According to elog14441, this box has flaky connectors which probably need to be remade
  2. Re-make the 64-pin IDC crimped connection on the cable from the coil driver board to sat. box, at the Satellite box end --- Chub and gautam

Any other ideas? The problem persists and it's annoying that the IMC cannot be locked.

  14841   Mon Aug 12 17:36:04 2019 gautamUpdateCDSMore bench test of c1iscaux

[chub, gautam]

With Chub's help, most of the problems have been resolved. Summary: I judge that we are good to go ahead with an install tomorrow.

  1. The problem with the BIO channels was a mis-wiring internal to the chassis - Chub fixed this and now all 32 AA enable/disable switches seem to work as advertised. Of course we will need to do the in-situ test to make sure.
  2. The problem with the ADC channels were multiple:
    • On the software end, I had gotten some addressing wrong - this was fixed.
    • On the hardware side - even though the inputs of the Acromag are "differential", I found that the readback was extremely noisy (~0.5 V RMS for a 3 V DC signal from the handheld calibrator unit 😲 ). Looking through the manual, I found a recommendation (pg10) that the "IN-" terminal of the Acromag ADC units be tied to the "RTN" pins on the same units. I don't know if this preserves the differential receiving capability of the Acromag ADCs - anyways, after Chub implemented this change, all the Analog Input channels behave as expected (I tested with a DC voltage and also a 200 mHz sine wave from a function generator).
    • Note that most of the Eurocard electronics we use are single-ended sending anyways.
    • What does this mean for the other Acromag ADCs (e.g. OSEM Shadow Sensor monitors) we have installed????? I saw no documentation in the elog/wiki.
  3. Binary input channel:
    • This is used by the "CM LIMIT" channel.
    • I found that I had to initialize a separate alias for the BIO3 unit, which acquires this signal, to use the modbus function "4" corresponding to "Read Input Registers" - c.f. the binary output modbus function 6, which is to "Write Single Register".
    • The fix for the mbbo channels is also likely to be along this lines - but I don't have the energy for that endavor right now.
  4. Testing of the physical mbboDirect bit channels using the Acromag Window utility
    • I can't get the mbboDirect EPICS record to work as expected, so I decided to use the native Acromag utility to test the functionality
    • First I released control of the acromags from the supermicro (stopped modbus)
    • There were several wiring errors - Chub had left for the day so I just fixed it myself.
    • The LED tester kit was used to check that the correct bits were flipped - they were.
  5. At the time of writing, the non-functional channels (in EPICS) are all related to the CM board:
    • C1:LSC-CM_LIMIT (binary input) tested later in the day, works okay...
    • C1:LSC-CM_REFL1_BITS (mbboDirect)
    • C1:LSC-CM_REFL2_BITS (mbboDirect)
    • C1:LSC-CM_AO_BITS (mbboDirect)
    • C1:LSC-CM_BOOST2_BITS (mbboDirect)

Since we don't immediately need the CM board, I say we push ahead with the install - at least that will restore the ability to lock PRMI / DRMI. Then we can debug these issues in situ - I'm certain the issue is related to the EPICS/Modbus setup and not the hardware because I verified the physical channel map using the Acromag windows utility.

Remaining Tasks:

  1. Install power supply cables at 1Y3
  2. Install supermicro and Acromag crates in 1Y3
  3. Migrate existing P1 connectors to P2 where applicable (Whitening boards)
  4. Connect Dsub-->P1 / P2 adaptors
  5. Run in-situ tests
Quote:

I bench tested the functionality of all the c1iscaux Acromag crate channels. Summary: we are not ready for a Monday install, much debugging remains.

Attachment 1: iscauxCheclist.pdf
iscauxCheclist.pdf
  14840   Sun Aug 11 11:47:42 2019 gautamUpdateCDSBench test of c1iscaux

I bench tested the functionality of all the c1iscaux Acromag crate channels. Summary: we are not ready for a Monday install, much debugging remains.

  1. DAC channels were tested using 4 ch oscilloscope and stepping the whitening gain sliders through their 15 gain settings
    • Response was satisfactory - the output changes between 0 - 5 V DC in 15 steps.
    • This analog voltage is converted to binary representation by an on-board ADC on the whitening boards. So we may have to tune the offset voltage and range to avoid accidental bit flipping due to the analog voltage of a particualr step falling close to the bit-flipping edge of the on-board ADC. This will require an in-situ test.
    • Test passed
  2. BIO output channels were tested using a DMM, and monitoring the resistance between the BIO pin and the RTN pin. In the "ON" state, the expected resistance is ~5 Mohm, and in the off state, it is ~3 ohms.
    • The AA filter switches on BIO1 unit do not show the expected behavior - @ Chub, please check the wiring.
    • All others (except the mbboDirect bits, see next bullet) were okay, including those for the CM board that are NOT part of the mbboDirect groups.
    • Test failed
  3. ADC channels were tested by driving a ~2Vpp 300mHz sine wave with a function generator, and looking at the corresponding EPICS channel with StripTool.
    • I found that all the ADC channels don't function as expected.
    • Part of the problem is due to incorrect formatting of the EPICS records in the db files, but I think the ADCs also need to be calibrated with the precision voltage source.
    • Why only ADCs require calibration and not the DACs????
    • Test failed
  4. mbboDirect BIO output test - I made a little LED breadboard tester kit to simultaneously monitor the status of these groups of binary outputs.
    • The LSB is toggled as expected when moving the gain slider along.
    • However, the other bits in the group are not toggled correctly.
    • I believe this is a problem with either (i) the way the EPICS record is configured to address the bits or (ii) the incorrect modbus datatype is used to initialize the ioc.
    • It will be helpful if someone can look into this and get the mbboDirect bits working, I don't really want to spend more time on this.
    • Test failed

I am leaving the crate powered (by bench supplies) in the office area so I have the option to work remotely on this.

  14839   Fri Aug 9 20:58:33 2019 JonUpdateElectronicsBorrowed Variac transformer

I borrowed an old-looking Variac variable transformer from the power supplies cabinet along the y-arm. It is currently in the TCS lab.

  14838   Fri Aug 9 16:37:39 2019 gautamUpdateALSMore EY table work

Summary:

  1. 220 uW / 600 uW (~36 % mode-matching) of IR light coupled into fiber at EY.
  2. Re-connected the RF chain from the beat mouth output on the PSL table to the DFD setup at 1Y2.
  3. A beat note was found between the PSL and EY beams using the BeatMouth.

Motivation:

We want to know that we can lock the interferometer with the ALS beat note being generated by beating IR pickoffs (rather than the vertex green transmission). The hope is also to make the ALS system good enough that we can transition the CARM offset directly to 0 after the DRMI is locked with arms held off resonance.

Details:

Attachment #1: Shows the layout. The realized MM is ~36 %. c.f. the 85% predicted by a la mode. It is difficult to optimize much more given the tight layout, and the fact that these fast lenses require the beam to be well centered on them. They are reasonably well aligned, but I don't want to futz around with the pointing into the doubling crystal. Consequently, I don't have much control over the pointing.

Attachment #2: Shows pictures of the fiber tips at both ends before/after cleaning. The tips are now much cleaner.

The BeatMouth NF1611 DC monitor reports ~580 mV with only the EY light incident on it. This corresponds to ~60 uW of light making it to the photodiode, which is only 25% of what we send in. This is commensurate with the BS loss + mating sleeve losses.

To find the beat between PSL and EY beams, I had to change the temperature control MEDM slider for the EY laser to -8355 cts (it was 225 cts). Need to check where this lies in the mode-hop scan by actually looking at the X-tal temperature on the front panel of the EY NPRO controller - we want to be at ~39.3 C on the EY X-tal, given the PSL X-tal temp of ~30.61 C. Just checked it, front panel reports 39.2C, so I think we're good.

Next steps:

  • Fix the IMC suspension
  • Measure the ALS noise for the Y arm
  • Determine if improvements need to be made to the IR beat setup (e.g. more power? better MM? etc etc).

EY enclosure was closed up and ETMY Oplev was re-enabled after my work. Some cleanup/stray beam dumping remains to be done, I will enlist Chub's help on Monday.

Attachment 1: IMG_7791.JPG
IMG_7791.JPG
Attachment 2: fiberCleaning.pdf
fiberCleaning.pdf
  14837   Fri Aug 9 08:59:04 2019 gautamUpdateCDSPrep for install of c1iscaux

[chub, gautam]

We scoped out the 1Y3 rack this morning to figure out what needs to be done hardware wise. We did not think about how to power the Acromag crate - the LSC rack electronics are all powered by linear supplies and not Sorensens, and the linear supplies are operating at pretty close to their maximum current-drive. The Acromag box draws ~3A of current from the 20 V supply, not sure what the current draw will be from the 15 V supply. Options:

  1. Since there are sorensens in 1Y2 and 1Y1, do we really care about installing another pair of switching supplies (+20 V DC and +15 V DC) in 1Y3?
    • Contingent on us having two spare Sorensens available in the lab. Chub has already located one.
  2. Use the Sorensens installed already in 1Y1. 
    • Probably the easiest and fastest option.
    • +15 V already available, we'd have to install a +20 V one (or if the +/-5 V or +12 V is unused, reconfigure for +20 V DC).
    • Can argue that "this doesn't make the situation any worse than it already is"
    • Will require the running of some long (~3 m) long cabling to bring the DC power to 1Y3 where it is required. 
  3. Get new linear supplies, and hook them up in parallel with the existing.
    • Need to wait for new linear supply to arrive
    • Probably expensive
    • Questionable benefit to electronics noise given the uncharacterized RF pickup situation at 1Y2

I'm going with option #2 unless anyone has strong objections.

  14836   Thu Aug 8 12:01:12 2019 gautamUpdateIOOMC1 suspension oddness

At ~1am PDT today, all the MC1 shadow sensor readbacks (fast CDS channels and Slow Acromag channels, latter not shown here) went to negative values. Of course a negative value makes no sense. After ~3 hours, they came back to positive values again. But since then, the shadow sensor RMS noise has been significantly higher in the >20 Hz band, and there are frequent glitches which kick the suspension. The IMC has been having trouble staying locked. I claim that this has to do with the Satellite box.

No action being taken now while I work on the ALS. In the past the problem has fixed itself.

Attachment 1: MC1_suspension.png
MC1_suspension.png
Attachment 2: MC1_suspension.pdf
MC1_suspension.pdf
  14835   Tue Aug 6 23:09:20 2019 gautamUpdateALSEY table work
  • Removed power monitoring PD (It was off anyways)
  • Installed Steering mirror and collimator in K6XS mount (fast axis = p-pol to best effort by eye)
  • Installed lens mounts in approx position
  • Cleaned fiber at EY and connected to the collimator
  • Coupled EY--->PSL and spare PSL-->EY fibers together at the PSL table to facilitate coupling.
  • tbc tomorrow...
Quote:

Couple IR light into fiber with good MM at EY

  14834   Tue Aug 6 16:44:50 2019 yehonathanUpdateLoss Measurement 

I grab 2 hours of the PD measurements using dlData_simple.ipynb in the misaligned state.

I get pretty much a normally distributed reading without drifts (Attachements 1 and 2).

The error in the reading is ~ 0.5%.

 

I am pretty sure this amount of noise is enough to explain the big noise in the Loss figure measurement.

 

The reason is that the loss formula is #(1-P_Locked/P_Misaligned+T1)-T2) where T1 and T2 are the transmissions of the ITM and ETM.

The average of the ratio P_Locked/P_Misaligned is ~ 1.01 for a loss figure of ~ 100ppm.

The standard deviation of the ratio is ~ 1% which is also the standard deviation of the expression in the brackets.

The average of this experssion however is ~ 0.01.

The reduction of the mean amplifies the error in the loss measurments by a factor of a few 10s!

Attachment 1: figure_1.png
figure_1.png
Attachment 2: figure_1-1.png
figure_1-1.png
  14833   Tue Aug 6 15:52:06 2019 gautamUpdateBHDPreliminary BHD calculations

Summary:

The requirement on the phase noise on the direct backscatter from the OMC back into the SRM is that it be less than 10^{-5} \, \mathrm{rad/\sqrt{Hz}} \approx 10^{-12} \, \mathrm{m/\sqrt{Hz}} @ 100 Hz, for a safety factor (arbitrarily chosen) of 10 (= 20dB below unsqueezed vacuum). Assuming 5 optics between the OMC and SRM which contribute incoherently for a factor of sqrt(5), and assuming a total of 1 ppm of the LO power to be backscattered, we need the suspensions to be moving < 5 \times 10^{-13} \, \mathrm{m/\sqrt{Hz}} @ 100 Hz. This seems possible to realize with single stage suspensions - I assume we get f^4 filtering from the pendulum at 100 Hz, and that there is an additional 80 dB attenuation (from the stack) of the assumed 1 micron/rtHz motion at 100 Hz, for an overall 160 dB attenutaiton, yielding 10^-14 m/rtHz at 100 Hz.

Details:

This is the same calculation as I had posted a couple of months ago (see elog that this is a reply to), except that Koji pointed out that the LO power is expected to dominate the (carrier) power incident on the OMC cavity(ies). So the more meaningful comparison to make is to have the x-axes of the plots denote the backscatter fraction, \epsilon_{\mathrm{BS}} rather than the LO power. One subtlety is that because the phase of the scattered field is random, the displacement-noise induced phase noise could show up in the amplitude quadrature. I think that in these quadrature field amplitude units, the RIN and phase noise are directly comparable but I might have missed a factor of 2*pi. But in the worst case, if all the phase noise shows up in the amplitude quadrature, we end up being only ~10dB below unsqueezed vacuum (for 1 ppm backscatter). 

For the requirement on the noise in the intensity quadrature - I think this is automatically satisfied because the RIN requirement on the incident LO field is in the mid 10^-9 1/rtHz regime.

Attachment 1: OMCbackscatter.pdf
OMCbackscatter.pdf
  14832   Tue Aug 6 14:55:23 2019 gautamUpdateCDSMaking Matlab R2015b the default

ML2013 is unable to open Simulink on any of the workstations. We decided to make the default version of Matlab R2015b (the default of the version of RCG we are using).

I commenced the procedure of the migration, starting with making a tagged commit of the current running simulink models. A local backup was also made, plus we have the usual chiara-based backup so I think we're in good hands.

Currently the branch and tag are protected - once we verify that everything works as expected post migration, I will open it up. I changed the directory structure of the models, need to confirm that the rtcds compilers don't have any hardcoded paths which may break due to my change.

The symlink to Matlab R2013 was deleted and a new symlink to R2015b was made. I activated the license using the Caltech campus license. Now running matlab from shell starts up R2015b yes. Simulink even works 😲 .

Attachment 1: ML2015b.png
ML2015b.png
  14831   Tue Aug 6 14:12:02 2019 yehonathanUpdateComputersmaking rossa great again

cdsutils is not working on rossa.

Import cdsutils produces this error:

In [2]: import cdsutils
---------------------------------------------------------------------------
OSError                                   Traceback (most recent call last)
<ipython-input-2-949babce8459> in <module>()
----> 1 import cdsutils

/ligo/apps/linux-x86_64/cdsutils-480/lib/python2.7/site-packages/cdsutils/__init__.py in <module>()
     53 
     54 try:
---> 55     import awg
     56 except ImportError:
     57     pass

/ligo/apps/linux-x86_64/cdsutils-480/lib/python2.7/site-packages/cdsutils/awg.py in <module>()
     30 """
     31 
---> 32 import sys, numpy, awgbase
     33 from time import sleep
     34 from threading import Thread, Event, Lock

/ligo/apps/linux-x86_64/cdsutils-480/lib/python2.7/site-packages/cdsutils/awgbase.py in <module>()
     17 libawg = CDLL('libawg.so')
     18 libtestpoint = CDLL('libtestpoint.so')
---> 19 libSIStr = CDLL('libSIStr.so')
     20 
     21 ####

/ligo/apps/anaconda/lib/python2.7/ctypes/__init__.pyc in __init__(self, name, mode, handle, use_errno, use_last_error)
    364 
    365         if handle is None:
--> 366             self._handle = _dlopen(self._name, mode)
    367         else:
    368             self._handle = handle

OSError: libSIStr.so: cannot open shared object file: No such file or directory

  14830   Mon Aug 5 17:36:04 2019 yehonathanUpdateLoss Measurement 

We check for unexpected drifts in the PD reading (clipping and such). We put a pickoff mirror where the PD used to be and place the PD at the edge of the table such that the beam is focused on it (see attachment).

The arms are completley misaligned. We note the time of start of measurement to be 1249086917.

Attachment 1: 20190805_171511.jpg
20190805_171511.jpg
  14829   Mon Aug 5 17:23:26 2019 gautamSummaryComputersWiFi Settings on asia

The VEA laptop asia was configured to be able to connect to too many WiFi networks - it was getting conflicted in its default position at the vertex and trying to hop between networks, for some reason trying to connect to networks that had poor signal strength. I deleted all options from the known networks except 40MARS. Now the network connection seems much more stable and reliable.

  Draft   Mon Aug 5 16:28:41 2019 yehonathanUpdateLoss Measurementwhat is going on with the loss measurements ?

We hypothesize that the systematic error in the loss measurement can come from the fact that the requirement on the alignment of the cavity mirrors is not stringent enough.

We repeat the loss measurement with 50 measurements. This time we change the thresholds for the error signals of the dither-align in the measureArmLoss.py file from 0.5 to 0.3.

We repeat the analysis done before:

We plot the reflected power of the two states on top of each other:

This  time it appears there was no drift. The histogram of the loss measurement:

The mean is 104ppm and the variation is 27%.

What I notice this time is that the PD readings in the aligned and misaligned states are anti-correlated. This is also true in the previous run (where there was drift) when looking in the short time scales. I plot several time series to demonstrate:

I wonder what can cause this behaviour.

  14827   Mon Aug 5 14:47:36 2019 yehonathanUpdateLoss Measurement 

Summary:

I analyze the 100 reps loss measurement of the Y arm using the AnalyzeLossData.ipynb notebook.

The mean of the measured loss is ~ 100ppm and the variation between the repititions is ~ 27%.

 

In Detail

In the real measurement the misaligned and locked states are repeatedly switched between each other. I plot the misaligned and locked PD readings seperately over time.

There seems to be a drift that is correlated between the two readings. This is probably a drift in the power after the MC2. To verify, I plot the ratio between those readings and find no apparent drift:

The variation in the ratio is less than 1%. The loss figure, computed to be 1 minus this ratio times a big number, give a much worse variation. I plot the histogram of the loss figure at each repitition (excluding extremely bad measurements):

The mean is ~ 100ppm. And the variation is ~ 27%.

  14826   Sun Aug 4 14:39:41 2019 gautamUpdateGeneralsome lab activity
  1. Unresponsive c1psl, c1iool0, c1auxey and c1iscaux VME crates were keyed.
  2. c1psl channels were burt-restored, did a burtrestore, and re-locked the PMC. Tweaked the pointing into the PMC on the PSL table to increase the PMC transmission from ~0.69 to ~0.71.
  3. Re-locked IMC. Ran WFS offset script to relieve the ~100 DAC counts (~10 urad) DC offset from the WFS servos to the IMC suspensions (a serious calibration of this into physical units should be made part of the planned 40m WFS activity). Now that I think about it, since we change the IMC alignment to match the input beam alignment, some post-IMC clipping could modulate the power incident on the ITMs, which is a source of error for the arm cavity loss determination using DC reflection. We need a better normalizing data stream than the IMC transmission.
  4. The IFO_OVEREVIEW medm screen was modified such that the threshold for the PMC transmitted beam to be visible was lowered from 0.7 to 0.6, so that now there is a continuous beam line from the NPRO to the PRM when the IMC is locked even when the PMC transmission degrades by 5% due to thermally driven pointing drifts on the PSL table.
  5. The wmctrl utility on pianosa wasn't working so well, I wasn't able to use my usual locking MEDM autoconfig scripts. Turned out to be due to a zombie MEDM window which I killed with xkill, now it is working okay again.
  6. The misaligned XARM was re-aligned and the loss measuring PDA520 at the AS port was removed from the beam path (mainly to avoid ADC saturations the fringing Michelson will cause).
  7. I noticed that the ETMX Oplev HeNe SUM level has degraded to ~50% of its power level from 200 days ago [Attachment #1], may need a new HeNe here soon. @Chub, do we have spare HeNes in stock?

I want to collect some data with the arms locked to investigate the possibility/usefullness of having seismic feedforward implemented for the arms (it is already known to help the IMC length and PRC angular stability at low frequencies). To facilitate diagnostics I modified the file /users/Templates/Seismic/Seismic_vs_TRXTRYandMC.xml to have the correct channel names in light of Lydia's channel name changes in 2016. Looking at the coherence data, the alignment of the cartesian coordinate system of the Seismometers at the ends and the global interferometer coordinate system can be improved.

I don't know if for the MISO filter design if there is any difference in using TRX/TRY as the target, or the arm length control signal.

Data collection started at 1249018179. I've setup a script running in a tmux shell to turn off the LSC enable in 2 hours.

Attachment 1: ETMX_OL.png
ETMX_OL.png
Attachment 2: Seismic_TRXTRYandMC_4Aug2019.pdf
Seismic_TRXTRYandMC_4Aug2019.pdf
  14825   Fri Aug 2 17:07:33 2019 yehonathan, gautamUpdateLoss Measurement 

We run a loss measurement on the Y arm with 50 repetitions.

  14824   Fri Aug 2 16:46:09 2019 KruthiUpdateCamerasClean up

I've put the analog camera back and disconnected the 151 unit GigE. But I ran out of time and wasn't able to replace the beamsplitter. I've put all the equipments back to the place where I took them from. The chopper and beam dump mount, that Koji had got me for the scatterometer, are kept outside, on the table I was working on earlier, in the control room. The camera lenses, additional GigEs, wedge beamsplitter, 1050nm LED and all related equipments are kept in the GigE box. This box was put back into CCD cameras' cabinet near the X arm.

Note: To clean stuff up, I had entered the lab around 9.30pm on Monday. This might have affected Yehonathan's loss measurement readings (until then around 57 readings had been recorded).

Sorry for the late update.

  14823   Fri Aug 2 11:37:38 2019 gautamUpdateALSEY IR ALS Assay

Summary:

I'd like to confirm that the IR ALS scheme will work for locking. The X-arm performance so far has been encouraging. I want to repeat the characterization for the Y arm. So I inspected the layout on the EY table, and made a list of characterization tasks. The current EY beam routing is difficult to work with, and it will definitely benefit from a re-do. However, I don't know how much time I want to spend re-doing it, so for a start, I will just try and couple some amount of light into a fiber and bring it to the PSL table, and see what noise performance I get.

Details:

Attachment #1: Photo of the current beam layout. The powers indicated were measured with the Ophir power meter.

  • I measure an SHG conversion efficiency of 0.87 %/W, which is considerably lower than the ~3.7%/W that is theoretically expected, and 1.5%/W that is realized at EX.
  • Of the 0.5 mW of green light that is generated, I measure ~0.375 mW at the viewport into the EY chamber. So there is ~25 % loss in the green beam path on the EY table. Seems high to me.
  • The previous solution of coupling IR light into the fiber realized at EY was to use the SHG leakage IR beam. While there isn't a measurement showing that this dirty beam is noisier than a cleaner pickoff, I'd like to adopt the solution used at EX, which is to use the leakage beam from the first steering mirror in the NPRO beam path. This will allow better mode-matching and polarization control of the beam being coupled into the fiber, which at least in principle, translates to less phase noise.
  • However - the beam layout at the EY table offers much less freedom to work with this idea than EX. A constraint is the clamp that secures the enclosure to the optical table, labelled in the photo. Further behind it, the green steering optics occupy all available space. A more comprehensive photo of the EY table can be found here.
  • Off the top of my head, I don't see any other good open spots on the EY table where we could couple IR light into the fiber. 
  • One other change I'd like to make is to replace the first steering mirror after the NPRO head, which is currently a Y1 HR mirror, with a R=99% BS. This will make it easier to control the amount of power coupled into the fiber, which is something we'd like.

Attachment #2: A candidate mode-matching solution, given the constraints outlined above. It isn't great, with only 85% modematching even theoretically possible. The lenses required are also pretty fast lenses. But I think it's the best possible without a complete overhaul of the EY layout. I'm still waiting for the lens kit to arrive, but as soon as they get here, I will start this work.

Characterization tasks:

  • Characterize SHG at EY [done 7/28]
  • Characterize gPDH at EY (loop TFs, improve MM, PDH discriminant, check the polarization)
  • Couple IR light into fiber with good MM at EY [done with 36% MM 8/9]
  • Clean fiber at EY, and at the PSL table [done 8/9]
  • Make the PSL + Y IR beat [done 8/9]
  • Noise budget
Attachment 1: IMG_7780.JPG
IMG_7780.JPG
Attachment 2: Ey_MM_20190802.pdf
Ey_MM_20190802.pdf
  14822   Thu Aug 1 13:55:34 2019 DuoBureaucracyEquipment loanGpib module taken to QIL lab

vanna --> QIL.

gautam 20190804: The GPIB module + power supply were returned to me by Duo ~5pm today at the 40m.

  14821   Wed Jul 31 17:57:35 2019 KojiUpdateBHDOMC cavity geometry

4 deg is not an optimized number optimized for criteria, but to keep the cavity short width to 0.1m. But the justification of 4deg is found in Section 3 and 4 of T1000276 on Page 4.

Quote:

Question for Koji: how is the aLIGO OMC angle of incidence of ~4 degrees chosen? Presumably we want it to be as small as possible to minimize astigmatism, and also, we want the geometric layout on the OMC breadboard to be easy to work with, but was there a quantitative metric? Koji points out that the backscatter is also expected to get worse with smaller angles of incidence.

  14820   Wed Jul 31 14:44:11 2019 gautamUpdateComputersSupermicro inventory

Chub brought the replacement Supermicro we ordered to the 40m today. I stored it at the SW entrance to the VEA, along with the other Supermicro. At the time of writing, we have, in hand, two (unused) Supermicro machines. One is meant for EY and the other is meant for c1psl/c1iool0. DDR3 RAM and 120 GB SSD drives have also been ordered, but have not yet arrived (I think, Chub, please correct me if I'm wrong).

Update 20190802: The DDR3 RAM and 120 GB SSD drives arrived, and are stored in the FE hardware cabinet along the east arm. So at the time of writing, we have 2 sets of (Supermicro + 120GB HD + 4GB RAM).

Quote:

We should ask Chub to reorder several more SuperMicro rackmount machines, SSD drives, and DRAM cards. Gautam has the list of parts from Johannes' last order.

  14819   Wed Jul 31 09:41:12 2019 gautamUpdateBHDOMC cavity geometry

Summary:

We need to determine the geometry (= round-trip length and RoC of curved mirrors) of the OMC cavities for the 40m BHD experiment. Sticking to the aLIGO design of a 4 mirror bowite cavity with 2 flat mirrors and 2 curved mirrors, with a ~4deg angle of incidence, we need to modify the parameters for the 40m slightly on account of our different modulation frequencies. I've setup some infrastructure to do this analytically - even if we end up doing this with Finesse, it is useful to have an analytic calculation to validate against (also not sure if Finesse can calculate HOMs up to order 20 in a reasonable time, I've only seen maxtem 8). 

Attachment #1: Heatmap of the OMC transmission for the following fields:

  • Carrier TEM00 is excluded, but HOMs up to m+n=20 included for both the horizontal and vertical modes of the cavity.
  • f1 and f2 upper and lower sidebands, up to m+n=20 HOMs for both the horizontal and vertical modes of the cavity, including TEM00.
  • Power law decay assumed for the HoM content incident on the OMC - this will need to be refined
  • The white region is where the cavity isn't geometrically stable.
  • Green dashed line indicates a possible operating point, white dashed line indicates the aLIGO OMC operating point. On the basis of this modeling, we would benefit from choosing a better operating point than the aLIGO OMC geometric parameters.

Algorithm:

  1. Compute the round-trip Gouy phase, \phi_{\mathrm{gouy}}, for the cavity.
  2. With the carrier TEM00 mode resonant, compute the round-trip propagation phase, \phi_{\mathrm{prop}} = \frac{2 \pi f_{\mathrm{offset}} L_{\mathrm{rt}}}{c}, and the round-trip Gouy phase, \phi_{\mathrm{G}} = (m+n)\phi_{\mathrm{gouy}} for the \mathrm{TEM}_{mn} mode of the field, with f_{\mathrm{offset}} specifying the offset from the carrier frequency (positive for the upper sideband, negative for the lower sideband). For the aLIGO cavity geometry, the 40m modulation sidebands acquire ~20% more propagation phase than the aLIGO modulation sidebands.
  3. Compute the OMC transmission for this round-trip phase (propagation + Gouy).
  4. Multiply the incident mode power (depending on the power law model assumed) by the cavity transmission.
  5. Sum all the fields.

Next steps:

  1. Refine the incident mode content (and power) assumption. Right now, I have not accounted for the fact that the f2 sideband is resonant inside the SRC while the f1 sideband is not. Can we somehow measure this for the 40m? I don't see an easy way as it's probably power dependent?
  2. Make plots for the projection along the slices indicated by the dashed lines - which HOMs are close to resonating? Might give us some insight.
  3. What is the requriement on transmitted power w.r.t. shot noise? i.e. the colorbar needs to be translated to dBVac.
  4. If we were being really fancy, we could simultaneously also optimize for the cavity finesse and angle of incidence as well.
  5. Question for Koji: how is the aLIGO OMC angle of incidence of ~4 degrees chosen? Presumably we want it to be as small as possible to minimize astigmatism, and also, we want the geometric layout on the OMC breadboard to be easy to work with, but was there a quantitative metric? Koji points out that the backscatter is also expected to get worse with smaller angles of incidence.

The code used for the ABCD matrix calcs have been uploaded to the BHD modeling GIT (but not the one for making this plot, yet, I need to clean it up a bit). Some design considerations have also been added to our laundry list on the 40m wiki.

Attachment 1: paramSpaceHeatMap.pdf
paramSpaceHeatMap.pdf
  14818   Tue Jul 30 20:11:12 2019 ranaSummaryIMCIMC ASC: thoughts and hopes

One of the biggest challenges in LIGO is reducing the alignment control noise. If you haven't worked on it for at least a few years, it probably seems like a trivial problem. But all versions of LIGO since 2001 have been limited by ASC noise below ~50 Hz.

I think the 40m IMC is a good testbed for us to try a few approaches towards mitigating this noise in LIGO. The following is a list of steps to take to get there:

  1. Using step responses and TF measurements, characterize the full existing system: SISO loop shapes, cross-couplings, and how diagnonal is the input and output matrices of the WFS. In principle, since we have 2 WFS in reflection and 1 DC QPD in the MC2 transmission, we should have full sensing of all angular DoFs.
  2. Check the correct operation of the WFS heads and the whole RF chain. We want the gains in the system to be such that either the shot noise or the RF electronics noise of the head is the limiting broadband noise in the system.
  3. Balancing the gains and phases of the demodulated signals is tricky, because we have no good reference. Should we use the JenneAM laser or the PSL beam?
  4. Estimate the coupling from the angular feedback signal to the IMC length noise using (1) sine wave injections for linear coupling, and (2) broadband noise for nonlinear coupling.
  5. We think the bilinear noise is due to the beam spot motion modulating the angle to length coupling as sensed by the laser beam. If this is true, we can increase the low frequency gain to minimize the beam spot motion (is this true?).
  6. By sinusoidally driving the mirror angles we can measure the instantaneous beam spot positions. We can then derive the matrix required to convert from our angular sensors (WFS + QPD) into beam spot motion. We should modify our IMC-WFS real-time model to give us DAQ channels which are beam spot estimators.
  7. Build a simulation of an IMC which has WFS, QPD, shot noise, and seismic noise.
  8. Use our optimal linear-feedback design tools to make Angular loops which minimize the bilinear noise coupling.
  9. Build a nonlinear controller (neural networks: dense + CNN) that outperforms the linear one by estimating the beam spot motion continuously and driving the cavity length to cancel the angle-to-length noise. 

I think that steps 1-6 are well within our existing experience, but we should do it anyway so as to reduce the IMC beam motion at low frequencies, and also to reduce the 10-100 Hz frequency noise as seen by the rest of the interferometer.

Steps 7-8 are medium hard, but we can get some help from the CSWG in tackling it.

Step is pretty tough, but I would like to try it and also get some help from MLWG and CSWG to address it.

  14817   Tue Jul 30 09:13:31 2019 gautamUpdatePSLc1psl keyed, Agilent setup cleared
  1. IMC would not lock. c1psl EPICS channels were unresponsive. I keyed the crate and went through the usual burtrestore/PMC-relocking dance.
  2. While at 1X2, I decided to take this opportunity to clean up the AG4395 setup that has been setup there unused for several weeks now.
    • Unplugged the active probe connected via BNC-T connector to the mixer IF output.
    • Noticed that the active probe (S/N 2850J01450) did not have it's power connection connected. According to the manual, this is bad. I don't know if the probe is damaged or not.
    • Moved the AG4395 cart out of the way so that there is a little more room around 1X1/1X2.
  14816   Mon Jul 29 19:08:55 2019 yehonathanUpdateLoss MeasurementReviving loss measurement by reflection

1. X arm is totally misaligned in order to measure the Y arm loss using the reflection method. Each measurement round consists of measuring the reflected power when the Y arm is aligned and when it is misaligned.

2. The measurement script used is /scripts/lossmap_scripts/armLoss/measureArmLoss.py. It generates a log file in the /logs folder specifying the alignment and misalignment times.

3. The data extraction script dlData.py processes the raw data in the log file and creates a hdf5 file in the /Data folder conataining the data of the measurement it self.

4. dlData.py labels the the aligned and misaligned datas incorrectly when the number of measurement is odd. I use only even number of measurements then.

5. In order to clip the chaotic transition between the aligned and misaligned states I use tDur attribute smaller than the actual sleep time used in the measurement script itself.

6. plotData.py (written by Gautam) and AnalyzeLossData.ipynb (written by me) can be used to calculate the loss and to plot some analyses (see https://nodus.ligo.caltech.edu:8081/40m/14568). They give roughly the same answer. The descripancy can be explained by the different modulation and ITM transmissions used.

7. I take a measurement of 8 repeatitions. I plot the measured reflected power alternating between the aligned and misaligned states. 

I find that the reflected power is repeatable to within 1%.

This is consistent with the transmission data plotted here which is also repeatable to within 1%.

8. I take an overnight measurement of 100 repeatitions.

  14815   Mon Jul 29 13:32:56 2019 gautamUpdateLoss MeasurementLoss measurement PD installed in AS path

[yehonathan, gautam]

  • we placed a PDA520 photodiode in the AS beampath, so AS110 and AS55 no longer see any light.
  • ITMX and ETMX were misaligned (since the plan is to measure the Y arm loss).
  • The PDA520 and MC2 transmission are currently going to the Y arm ALS beat channels in the DAQ system. Unfortunately, we have no control over the whitening gains for these channels because of the c1iscaux2 situation.
  14814   Fri Jul 26 19:53:53 2019 JonOmnistructureCamerasGigE Camera Server

I've started setting up the last new rackmount SuperMicro as a dedicated server for the GigE cameras. The new machine is currently sitting on the end of the electronics test bench. It is assigned the hostname c1cam at IP 192.168.113.116 on the martian network. I've installed Debian 10, which will be officially supported until July 2024.

I've added the /cvs/cds NFS mount and plan to house all the client/server code on this network disk. Any dependencies that must be built from source will be put on the network disk as well. Any dependencies that can be gotten through the package manager, however, will be installed locally but in an automated way using a reqs file.

We should ask Chub to reorder several more SuperMicro rackmount machines, SSD drives, and DRAM cards. Gautam has the list of parts from Johannes' last order.

  14813   Thu Jul 25 20:08:36 2019 gautamUpdateComputersSolidworks machine

I brought one CPU (Dell T3500) and one 28" monitor from Mike Pedraza's office in Downs to the 40m. It is on Steve's desk right now, pending setup. The machine already has Solidworks and Altium installed on it, so we can set it up at our leisure. The login credentials are pasted on the CPU with a post-it should anyone wish to set it up.

  14812   Thu Jul 25 14:28:03 2019 gautamConfigurationComputersfirewalld disabled for EPICS CA

I think rana did some more changes to this workstation to make it useful for commissioning activities - but the MEDM screens were still white blanks. The problem was that the firewalld wasn't disabled (last two steps of the KThorne setup wiki). I disabled it. Now donatella can run MEDM, ndscope and StripTool. DTT doesn't work to get online data because of a "Synchronization Error", I'm not bothering with this for now. I think Kruthi successfully demonstrated the fetching of offline data with DTT.

Attachment 1: donatellaCommissioning.png
donatellaCommissioning.png
  14811   Thu Jul 25 12:25:56 2019 gautamUpdateALSIR ALSX noise

Summary:

  1. There are some broad peaks in the ALS out-of-loop noise, centered at ~145 Hz, ~245 Hz and ~570 Hz which are absent in both the POX in-lock error signal and in the green PDH error signals (see Attachment #1). So I conclude they originate in the IR ALS beat chain somewhere. Needs more investigation, in the general quest to improve the ALS noise.
  2. This measurement also shows that the ALS noise is limited by unsuppressed EX green PDH frequency noise above ~400 Hz (100 Hz if you ignore the unexplained broad humps).

These spectra were taken with the arm cavity length locked to the PSL frequency using POX as an error signal, and the EX laser frequency locked to the XARM cavity length by the analog PDH servo at EX, so there is no feedback control with the ALS beat signal as an error signal.

Other details:

  • The transition of arm resonance control from POX to ALS error signal is more robust now - I am able to do this during daytime, and also maintain the lock for >20 minutes at a time.
  • Rana encouraged me not to spend too much time on this - so my next goal here will be to get the Y arm IR ALS working, and then we can control the two arms using ALS error signals in the CARM/DARM basis instead of the X/Y basis.
  • I still think it's worth getting the ALS good enough that the locking becomes repeatable and reliable.
    • The main task here is going to be re-doing the EY green layout to match the EX layout, get good MM into the cavity etc.
    • The IR light also has to be coupled into the fiber at EY.
Attachment 1: ALS_broadPeaks.pdf
ALS_broadPeaks.pdf
  14810   Thu Jul 25 09:19:32 2019 JonUpdateCamerasUpgraded Pylon from 5.0.12 to 5.2.0

I'll keep developing the camera server on a parallel track using the "new_..." directory naming convention. One thing I forgot to note is that the new pylon/pypylon packages require Python 3, so will not work with any of the 2.7 scripts. All of the environment I need to set up is exclusively Python 3. I won't change anything in the Python 2.7 environment in current use.

Also, I found the source of the bit resolution issue: Joe B's code loads a set of initialization parameters from a config file. One of them is "Frame Type = Mono8" which sets the dynamic range of the stream. I'll look into how this should be changed. 

Quote:

Since there are multiple SURF projects that rely on the cameras:

  1. I moved the new installs Jon made to "new_pylon5" and "new_pypylon". The old installs were moved back to be the default directories.
  2. The bashrc alias for pylon was updated to allow the recording of videos (i.e. it calls the PylonViewerApp from new_pypylon).
  3. There is a script that can grab images at multiple exposures and save 12-bit data as uint16 numpy arrays to an HDF5 file. Right now, it is located at /users/kruthi/scripts/grabHDR.py. We can move this to a better place later, and also improve the script for auto adjusting the exposure time to avoid saturations.
  14809   Thu Jul 25 00:26:47 2019 MilindUpdateCamerasConvolutional neural networks for beam tracking

Somehow I never got around to doing the pixel sum thing for the new real data from the GigE. Since I have to do it for the presentation, I'm putting up the results here anyway. I've normalized this and computed the SNR with the true readings.

SNR = (power in true readings)/ (power in error signal between true and predicted values)

Attachment #2 is SNR of best performing CNN for comparison.

Attachment 1: centroid.pdf
centroid.pdf
Attachment 2: subplot_yaw_test.pdf
subplot_yaw_test.pdf
  14808   Wed Jul 24 20:23:52 2019 gautamUpdateCamerasUpgraded Pylon from 5.0.12 to 5.2.0

Since there are multiple SURF projects that rely on the cameras:

  1. I moved the new installs Jon made to "new_pylon5" and "new_pypylon". The old installs were moved back to be the default directories.
  2. The bashrc alias for pylon was updated to allow the recording of videos (i.e. it calls the PylonViewerApp from new_pypylon).
  3. There is a script that can grab images at multiple exposures and save 12-bit data as uint16 numpy arrays to an HDF5 file. Right now, it is located at /users/kruthi/scripts/grabHDR.py. We can move this to a better place later, and also improve the script for auto adjusting the exposure time to avoid saturations.

My changes were necessary because the grabHDR.py script was throwing python exceptions, whereas it was running just fine before Jon's changes. We can move the "new_*" dirs to the default once the SURFs are gone.

Let's freeze the camera software config in this state until next week.

  14807   Wed Jul 24 20:05:47 2019 MilindUpdateCamerasCNNs for beam tracking || Tales of desperation

At the lab meeting today, Rana suggested that I use the Pylon app to collect more data if that's what I need. Following this, Jon helped me out by updating the pylon version and installing additional software to record video. Now I am collecting data at

  1. higher exposure rate - 600 us magically gives me a saturation percentage of around 1%, see attachment #1 (i.e around 1% of the pixels in the region containing the beam spot are over 240 in value). Ths is a consequence of my discussion with Gabriele where we concluded that I was losing information due the low exposure rate I was using.
  2. For much longer: roughly 10 minutes
    1. at an amplitude of 40 cts for 0.2 Hz
    2. at an amplitude of 20 cts for 0.2 Hz
    3. at an amplitude of 10 cts for 0.2 Hz
    4. at an amplitude of 40 cts for 0.4 Hz
    5. at an amplitude of 20 cts for 0.2 Hz
    6. Random motion

Consequently I have dithered the MC2 optic from around 9:00 PM.

Attachment 1: saturation_percentage.pdf
saturation_percentage.pdf
  14806   Wed Jul 24 16:45:32 2019 JonUpdateCamerasUpgraded Pylon from 5.0.12 to 5.2.0

I upgraded Pylon, the C/C++ API for the GigE cameras, to the latest release, 5.2.0. It is installed in the same location as before, /opt/rtcds/caltech/c1/scripts/GigE/pylon5, so environment variables do not change. The old version, 5.0.12, still exists at opt/rtcds/caltech/c1/scripts/GigE/backup_pylon5.

The package contains a GUI application (/bin/PylonViewerApp) for streaming video. The old version supports saving still images, but Milind discovered that the new version supports saving video as well. This required installing a supplementary package supporting MPEG-4 output.

Basler's GUI application is launched from the terminal using the alias pylon. I've tested it and confirm it can save both videos and still-image formats. I recommend to also try grabbing images using this program and check the bit resolution. It would be a useful diagnostic to know whether it's a bug in Joe B.'s code or something deeper in the camera settings.

Attachment 1: IMG_3525.jpg
IMG_3525.jpg
  14805   Wed Jul 24 12:24:43 2019 MilindUpdateIOOunstick.py and ifotest

Moved the unstick.py code to the ifotest repository here. It now handles signals like those generated by Ctrl-C and so forth. It can still be run as python unstick.py <machine1> <machine2> etc.

  14804   Wed Jul 24 04:20:35 2019 KruthiUpdateCalibration-RepairMC2 pitch and yaw calibration

Summary:  I calibrated MC2 pitch and yaw offsets to spot position in mm. Here's what I did:

  1. Changed the MC2 pitch and yaw offset values using  ezca.Ezca().write('IOO-MC2_TRANS_PIT_OFFSET', <pitch offset value> ) and ezca.Ezca().write('IOO-MC2_TRANS_YAW_OFFSET', <yaw offset value> )
  2. Waited for ~ 700-800 sec for system to adjust to the assigned values
  3. Took snapshots with the 2 GigEs I had installed - zoomed in and zoomed out. (I'll be using these to make a scatter loss map, verify the calibration results, etc)
  4. Ran the mcassDecenter script, which can be found in /scripts/ASS/MC. This enters the spot position in mm in the specified text file.

Results:  In the pitch/yaw vs pitch_offset/yaw_offset graph attached,

  • intercept_pitch = 6.63 (in mm) ,  slope_pitch = -0.6055 (mm/counts) 
  • intercept_yaw = -4.12 (in mm) ,  slope_yaw = 4.958 (mm/counts) 
Attachment 1: Pitchyaw_calibration.png
Pitchyaw_calibration.png
ELOG V3.1.3-